# Gyrokinetic flux-driven simulations in mixed TEM/ITG regime using a delta-f PIC scheme with evolving background

### v1

@@ I Introduction
Magnetic fusion research relies heavily on its accurate modeling by computer simulations. In the most promising reactor configuration, the tokamak, the plasma is confined by magnetic fields in a toroidal vacuum chamber. A complete description of the plasma involves simulating regions of the core, edge, the Scrap-Off-Layer (SOL), and plasma-wall interaction.
@@
Following the success of a previous work Murugappan et al. (2022) by the author using a similar approach in a simplified setup (sheared slab geometry, adiabatic electrons), the adaptive scheme is implemented in tokamak relevant axisymmetric toroidal geometry in the frame of the global gyrokinetic PIC code ORB5 Lanti et al. (2020). The simulations are electrostatic and ‘flux-driven’, with fully kinetic ions while electrons have a hybrid response Lanti et al. (2018) and instabilities being driven by both ions and electrons in a mixed Ion-Temperature-Gradient-Trapped-Electron-Mode (ITG-TEM) regime. Background density and temperature adaptation are made for both species independently. When compared to non-adaptive cases, results from the adaptive scheme exhibit low errors resulting from statistical sampling noise. Evolving the plasma profiles in the presence of radially localized sources up to their quasi-steady state using the adaptive scheme will be shown to require much lower marker numbers than with the non-adaptive scheme.
@@
This paper is organized as follows. Sec. II introduces the physical model to be solved. Namely, the Vlasov-Maxwell equation in the first order gyrokinetic approximation. A discussion of the Quasi-Neutrality-Equation (QNE) then follows, elaborating on the electrons’ hybrid response. It concludes by describing the functional form of the different possible source terms used in this work. Sec. III elaborates on the use of a control variate in the form of either a canonical or local Maxwellian, with flux-surface-averaged (f.s.a.) density, parallel flow and temperature having an explicit time-dependence. The relaxation equations that connect the three lowest order velocity moments of $f_{0}$ and $\delta f$ are then explained. The modification of $\delta f$ requires a correction term to be added to the right-hand-side (r.h.s.) of the QNE. Sec. IV details the considered initial profiles of density and temperature, along with the form of the actual heat sources for each species used in this work. Information about grid resolution, Fourier filtering, source strengths, and adaptive scheme parameters are given. Sec. V.1 begins the result section with a discussion on time traces of various transport parameters and sampling noise diagnostics, comparing them between non-adaptive and adaptive cases. Sec. V.2 investigates the effect of the adaptive scheme on time-averaged profiles of f.s.a. density and temperature. This section also examines the need of the QNE r.h.s. correction. Sec. V.3 diagnoses the time evolution of sampled phase-space volume and marker distribution, which reveals in some simulation cases a problem of under-sampling. Sec. V.4 discusses the f.s.a. profiles of density and temperature, time-averaged over te quasi-steady state, results of which could acquired only under the adaptive scheme. Sec. VI then summarizes the main merits of the adaptive scheme demonstrated in this work, along with possible future developments and applications directions.
@@ II Physical model
where $S_{\rm j}$ is a general source term. Here, $f_{\rm j}$ is the gyrocenter distribution in $5$D gyrocenter phase space $\vec{Z}=[\vec{R},v_{\parallel},\mu]$, where $\vec{R}$ is the $3$D configuration space vector, $v_{\parallel}$ the parallel velocity and $\mu=v_{\perp}^{2}/2B$ the magnetic moment per mass, with $v_{\perp}$ the gyration velocity and $B$ the strength of the local magnetic field $\vec{B}=B\hat{b}$. ORB5 uses magnetic coordinates $[s,{\theta^{\star}},\varphi]$ for representing the gyrocenter position $\vec{R}$ in tokamak geometry, with $s=\sqrt{\psi/\psi_{\rm edge}}$ the normalized radial coordinate expressed in terms of the poloidal magnetic flux function $\psi$ and its value at the edge of the radial boundary $\psi_{\rm edge}$. ${\theta^{\star}}$ is the straight-field-line poloidal angle, and $\varphi$ is the toroidal angle. As this work is concerned with electrostatic turbulence, the equations of motion of $\vec{Z}$ are given by
@@
In order to reach quasi-steady state, it was shown Krommes (1999) that some form of dissipation is required. Given that we neglect physical collisions, we therefore introduce a noise control Krook operator $S_{\rm nj}=-\gamma_{\rm n}(f_{\rm j}-f_{\rm 0j})+S_{\rm nj}^{\rm c}$ of rate $\gamma_{\rm n}$, which is species independent and is taken to be a small fraction (typically around $10\%$) of the maximum linear growth rate of the instabilities driving the turbulence. $S_{\rm nj}$ relaxes the distribution $f_{\rm j}$ to the background $f_{\rm 0j}$, while conserving f.s.a. density, parallel moment, residual zonal flows, and energy. These conservation properties are ensured by the correction term $S_{\rm nj}^{\rm c}$. Besides serving as a noise control operator for the PIC scheme, temperature-gradient-driven simulations of this work also uses the source term $S_{\rm j}$ as a fixed power heat source (without f.s.a. energy conservation). For flux-driven simulations, the f.s.a. heating operator used is given by
@@
Here, $\mathcal{E}=v^{2}/2=v_{\parallel}^{2}/2+\mu B$ is the kinetic energy per mass of the gyrocenter. The heat source $S_{\rm H}$ is parameterized by the f.s.a. profiles $n_{\rm H}$ and $T_{\rm H}$ appearing in Eq.(8), which in this work are taken to be the initial $n_{0}$ and $T_{0}$ profiles of each species. $\gamma_{\rm H}$ and $G_{\rm H}$ of Eq.(7) represent the heating rate and radial heating profile normalized to the unit of temperature, respectively. The form $\frac{\partial f_{\rm L}}{\partial T_{\rm H}}$ analytically ensures no f.s.a. density or momentum source. Nonetheless, $S_{\rm H}^{\rm c}$ ensures numerical conservation of f.s.a. density, parallel momentum, and residual zonal flows to round-off. Finally, to damp turbulence at the outer radial edge ($s=1$) of the simulation domain, a buffer in the form of a Krook operator $S_{\rm b}(s)=-\gamma_{\rm b}[(s-s_{\rm b})/(1-s_{b})]^{4}\delta f$, of species-independent damping rate $\gamma_{\rm b}$ is used. $S_{\rm b}$ is parametrized by the species-dependent buffer entrance $s=s_{\rm b}$ and only acts in the radial range $s\in[s_{\rm b},1]$. Taken together, the r.h.s. of Eq.(1) now reads
@@ III Numerical methods
Under the delta-f splitting defined by Eq.(4), it is common practice to choose $f_{0}=f_{\rm eq}$, where $f_{\rm eq}$ is an equilibrium distribution of the unperturbed system in the absence of the heat source term $S_{\rm H}$. This implies that $f_{\rm eq}$ has to be a function of the constants of motion of the unperturbed system. Namely, the modified canonical toroidal momentum Angelino et al. (2006) ${\hat{\psi}_{0}}=\psi+F(\psi)v_{\parallel}/\Omega_{\rm c}+\psi_{0corr}$, the kinetic energy $\mathcal{E}$, and the magnetic moment $\mu$. Here, $F(\psi)$ is the poloidal current flux function, as it appears in the relation for the axisymmetric equilibrium magnetic field $\vec{B}=F(\psi)\nabla\varphi+\nabla\psi\times\nabla\varphi$, and $\psi_{0corr}$ is a correction term that results in ${\hat{\psi}_{0}}\sim\psi$ on average along unperturbed trajectories. Under the Monte-Carlo interpretation Aydemir (1994), $f_{0}$ serves as a control variate, provided that the delta-f assumption
@@
Note that time dependence enters the control variate of Eq.(LABEL:eq:f0lm) via the f.s.a. background profiles of density $n_{0}$, parallel flow $u_{0}$ and temperature $T_{0}$ of the gyrocenters. It is the form of Eq.(LABEL:eq:f0lm) which will be used to develop the time dependent control variate scheme to ensure that Eq.(9) remains valid throughout simulations.
@@
must be appended. Given the analytic form for $f_{0}$ from Eq.(LABEL:eq:f0cm) or (LABEL:eq:f0lm), Eq.(LABEL:eq:qne_corr) is calculated using quadratures on a 5D phase space grid. An alternative, and computationally cheaper way to avoid appending the correction term is to equate the two terms of Eq.(LABEL:eq:qne_corr). This means that the change in the background electron density is set to be approximately equal to the change in the background ion gyrodensity, i.e.
@@ IV Profiles and simulation parameters
where $\rho_{V}$ is the radial coordinate $\rho_{V}(\psi)=\sqrt{V(\psi)/V(\psi_{\rm edge})}$, with $V(\psi)$ the volume enclosed by the flux surface label $\psi$. $g_{0}$ and $g_{2}$ are coefficients determined such that $G$ and $\,\text{d}G/\,\text{d}\rho_{V}$ are continuous at $\rho_{V}=\rho_{\rm core}$, and $G_{\rm ped}=G_{1}+\mu_{G}(1-\rho_{\rm ped})$. The functional form, was found to describe well the experimentally observed profiles in TCV L-mode discharges Sauter et al. (2014). The values of the parameters in Eq.(LABEL:eq:profile) for the ion and electron density and temperature profiles are shown in Tab. 1. All densities and temperatures are normalized by $\bar{n}=N_{\rm ph}/V(\psi_{\rm edge})$ (see Eq.(18) for the definition of $N_{\rm ph}$) and $T_{0e}(s_{0})$ respectively.
@@
The grid resolution for the radial $s$, poloidal ${\theta^{\star}}$ and toroidal $\varphi$ is taken to be $N_{s}\times N_{{\theta^{\star}}}\times N_{\varphi}=256\times 512\times 256$, where $N$ represents the number of intervals. Toroidal Fourier modes in the range of $0\leq n\leq 64$ will be simulated, only keeping poloidal Fourier modes $m\in[nq(s)-\Delta m,nq(s)+\Delta m]$ with half-width $\Delta m=5$. TO lighten the computational cost ‘heavy’ electrons are considered, i.e. with an electron-ion mass ratio $m_{e}/m_{i}=1/200$. The time resolution is $c_{s}\Delta t/a=4.1\times 10^{-3}$, and the maximum linear growth rate is found to be $\gamma_{\rm max}a/c_{s}=0.931$. The strength of the Krook operator $S_{\rm n}$ for noise control is set to $\gamma_{\rm n}=11\%\gamma_{\rm max}$. For the flux-driven simulations considered here, the radial heat source profile $G_{\rm H}(\psi)$ of Eq.(7) is approximated by fitting a Gaussian function around the peak of the time-averaged effective heat source at quasi-steady state of a previously run temperature-gradient-driven simulation with the initial profiles given by Tab. 1, as the fitting of a Gaussian function best represents heat deposition profiles in experiments. The heat source strength $\gamma_{\rm H}$ is determined by equating the integrated power to that of the effective heat source. The heat sink at the radial edge is replaced by the Krook buffer $S_{b}$. The calculation for the heat source is done for both ions and electrons separately. An example of the heat source profiles for ions and electrons are shown in Fig.2.
Simulations are initialized by loading markers that represent $\delta f(t=0)$ as white noise. All adaptive cases have adaptation rates $\alpha_{n}=\alpha_{E}=11\%\gamma_{\rm max}$ and adaptation period set with $N_{t}=100$ for both ions and electrons. Even when considering the hybrid electron model, adapted electron profiles include contributions from both passing and trapped electrons. Unless stated otherwise, all time traces have a moving time averaging window of $c_{s}\Delta t/a=4.085$. Marker numbers $N_{p}$ are displayed in millions (M). Comparison will be made between the non-adaptive (standard) case with $N_{p}=256$M, $128$M and the adaptive case with $N_{p}=128$M, $64$M. The choice of simulations with different $N_{p}$ allows for the discerning of sampling noise’s effect on different diagnostic measures.
(a) Ion
(b) Electron
@@ V.1 Time traces
We get an overview of transport time scales by first considering time traces of heat diffusivity $\chi$, which are derived from the heat flux $\vec{q}_{\rm H}$ resulting from gyrocenter $E\times B$ drifts. $\vec{q}_{\rm H}$ is composed of the contributions from the kinetic energy $\vec{q}_{\rm kin}$, the field potential $\vec{q}_{\rm pot}$, and the particle flux $\vec{\Gamma}$, all of which are given respectively by Villard et al. (2013)
@@
The heat diffusivity $\chi$ is then given by
@@
where $n$ and $T$ are the gyrocenter density and temperature respectively. The diffusivity is usually given in gyroBohm units $\chi_{GB0}=\rho_{s}^{2}c_{s}/a$ evaluated at $s=s_{0}=1$. Figure 3 shows the ion heat diffusivity $\chi_{i}$ for all cases. Within the standard (non-adaptive) and adaptive cases, results appear converged in marker number $N_{p}$. The ion heat $\chi_{i}$ values of the adaptive cases are shown to be consistently higher than the standard cases. This is supported by the fact that, for a given flux level, the adaptive cases exhibit lower temperature gradients in $s\in[0.7,0.9]$ (see Fig.9(a)) as compared to the standard case. This difference between results of the two schemes can be explained by considering the shearing rate. Figure 4 shows the $s\in[0.7,0.9]$ radially averaged absolute value of the turbulence driven $E\times B$ zonal flow shearing rate, which is defined by Villard et al. (2002)
@@
To see how electron diffusivity $\chi_{e}$ compares to $\chi_{i}$, Fig.3(b) shows the ratio $\chi_{i}/\chi_{e}$ for all cases considered. During $c_{s}t/a\in[0,100]$, the edge region ($s\in[0.7,0.9]$) with almost constant gradient (see Fig.10(a)) is predominantly driven by TEM i.e. $\chi_{i}/\chi_{e}<1$. From $c_{s}t/a=200$ onwards, the system is in a mixed regime of instabilities driven by TEM and Ion-Temperature-Gradient (ITG) mode reflected by $\chi_{i}/\chi_{e}\sim 1$. This change of regime is due essentially to the density profile evolution, as will be discussed further (see Fig.11).
@@
As the system approaches quasi-steady state, with the buffer being the only source/sink of particles, particle flux is expected to reduce and reach near zero values. Figure 5 shows the f.s.a. ion gyrocenter flux $\left\langle\vec{\Gamma}\cdot\nabla\psi\right\rangle/\left\langle|\nabla\psi|\right\rangle$ for cases with various $N_{p}$ under standard and adaptive schemes. We first note that adaptive cases (blue and green curves) have consistent slightly higher ion gyrocenter flux compared to the standard cases (black and orange) right after the initial overshoot in the time window $c_{s}t/a\in[50,200]$. This is again consistent with the lower $\omega_{E\times B}$ amplitudes of the adaptive cases (see Fig.4), thereby allowing for higher turbulence levels, and therefore, particle transport to develop. Nonetheless, the standard case with $N_{p}=256$M (black line) seems to initially converge to the same quasi-steady state value as that of the adaptive cases.
@@
with $\left\langle w\right\rangle(s)=\sum_{p\in{\rm bin}}w_{p}/N_{\rm bin}$ and $\left\langle w^{2}\right\rangle(s)=\sum_{p\in{\rm bin}}w_{p}^{2}/N_{\rm bin}$ the expectation value of $w$ and $w^{2}$ within that bin, respectively. Here, $N_{\rm bin}$ is the number of markers belonging to the bin. Note that, one expects convergence of $\sigma_{w}$ with marker number $N_{p}$ at a given time $t$. Furthermore, with dissipation in the form of the Krook operator $S_{n}$, the growth of $\sigma_{w}$ will be limited at quasi-steady state. The measure of weight root-mean-square average $\sqrt{\left\langle w^{2}\right\rangle}$ is not used here as $\sigma_{w}$ accounts for the generally non-zero density perturbation i.e. $\left\langle w\right\rangle^{2}>0$. Figure 6 shows $\sigma_{w}$ for both standard and adaptive cases with marker number $N_{p}=128$M, up to a simulation time of $c_{s}t/a=600$. From Figs.6(a) and 6(b), we see that the radial location of increasing maximum values of $\sigma_{w}$ for both the ions and electrons are located where the profile deviation is the greatest (see Figs.10 and 11). After an initial phase where $\sigma_{w}$ grows (however to significantly lower values than in standard cases) the corresponding results under the adaptive scheme, on the other hand, have diminishing maxima. This is due to the adaptive control variate $f_{0}(t)$, with adaptive density and temperature profiles. Furthermore for the adaptive cases, we see a drop in electron $\sigma_{w}$ around $c_{s}t/a=50$, compared to that of the ion which drops gradually after $c_{s}t/a\simeq 200$, where $\sigma_{w}$ is gradually decreased. This can be explained by considering Fig.3(b), where we see that for $c_{s}t/a\in[0,50]$, the electron diffusivity $\chi_{e}$ is more than $200\%$ that of the ion $\chi_{i}$. This suggests a quicker evolving electron temperature profile $T_{e}$ during this period. Thus, as the adaptation rate is sufficiently high for this case, the electron $\sigma_{w}$ is reduced at a shorter time scale. Comparing results between the standard and adaptive cases, we see that the latter only have $15\%$ the maximum value of $\sigma_{w}$ compared to that of the former. Furthermore, at $c_{s}t/a=600$, $\sigma_{w}$ for the standard cases is still growing, while that of the adaptive cases exhibit a steady value of around $0.2$.
(a) Standard, ions
(b) Standard, electrons
(c) Adaptive, ions
(d) Adaptive, electrons
@@
Figure 8: Evolution up to time $c_{s}t/a=123$ of the radial profile of the zonal flow $E\times B$ shearing rate $\omega_{E\times B}$ with marker number $N_{p}=128$M. (b) has Eq.(LABEL:eq:qne_corr) calculated via quadratures, while (c) employs Eq.(LABEL:eq:qnrhs_fsa). All plots share the same color scale.

### v3

@@ I Introduction
Magnetic fusion research relies heavily on its accurate modeling by computer simulations. In the most promising reactor configuration, the tokamak, the plasma is confined by magnetic fields in a toroidal vacuum chamber. A complete description of the plasma involves simulating regions of the core, edge, the Scrape-Off-Layer (SOL), and plasma-wall interaction.
@@
Following the success of a previous work Murugappan et al. (2022) by the author using a similar approach in a simplified setup (sheared slab geometry, adiabatic electrons), the adaptive scheme is implemented in tokamak relevant axisymmetric toroidal geometry in the frame of the global gyrokinetic PIC code ORB5 Lanti et al. (2020). The simulations are electrostatic and ‘flux-driven’, with gyrokinetic ions while electrons have a hybrid response Lanti et al. (2018) and instabilities being driven by both ions and electrons in a mixed Ion-Temperature-Gradient-Trapped-Electron-Mode (ITG-TEM) regime. Background density and temperature adaptations are made for both species independently. When compared to non-adaptive cases, results from the adaptive scheme exhibit low errors resulting from statistical sampling noise. Evolving the plasma profiles in the presence of radially localized sources up to their quasi-steady state using the adaptive scheme will be shown to require much lower marker numbers than with the non-adaptive scheme.
@@
This paper is organized as follows. Sec. II introduces the physical model to be solved. Namely, the Vlasov-Maxwell equation in the first order gyrokinetic approximation. A discussion of the Quasi-Neutrality-Equation (QNE) then follows, elaborating on the electrons’ hybrid response. It concludes by describing the functional form of the different possible source terms used in this work. Sec. III elaborates on the use of a control variate in the form of either a canonical or local Maxwellian, with flux-surface-averaged (f.s.a.) density, parallel flow and temperature having an explicit time-dependence. The relaxation equations that connect the three lowest order velocity moments of $f_{0}$ and $\delta f$ are then explained. The modification of $\delta f$ requires a correction term to be added to the right-hand-side (r.h.s.) of the QNE. Sec. IV details the considered initial profiles of density and temperature, along with the form of the actual heat sources for each species used in this work. Information about grid resolution, Fourier filtering, source strengths, and adaptive scheme parameters are given. Sec. V.1 begins the result section with a discussion on time traces of various transport parameters and sampling noise diagnostics, comparing them between non-adaptive and adaptive cases. Sec. V.2 investigates the effect of the adaptive scheme on time-averaged profiles of f.s.a. density and temperature. This section also examines the need of the QNE r.h.s. correction. Sec. V.3 diagnoses the time evolution of sampled phase-space volume and marker distribution, which reveals in some simulation cases a problem of under-sampling. Sec. V.4 discusses the f.s.a. profiles of density and temperature, time-averaged over the quasi-steady state, results of which could be acquired only under the adaptive scheme. Sec. VI then summarizes the main merits of the adaptive scheme demonstrated in this work, along with possible future developments and applications directions.
@@ II Physical model
where $S_{\rm j}$ is a general source term. Here, $f_{\rm j}$ is the gyrocenter distribution in $5$D gyrocenter phase space $\vec{Z}=[\vec{R},v_{\parallel},\mu]$, where $\vec{R}$ is the $3$D configuration space vector, $v_{\parallel}$ the parallel velocity and $\mu=v_{\perp}^{2}/2B$ the magnetic moment per mass, with $v_{\perp}$ the perpendicular velocity and $B$ the strength of the local magnetic field $\vec{B}=B\hat{b}$. ORB5 uses magnetic coordinates $[s,{\theta^{\star}},\varphi]$ for representing the gyrocenter position $\vec{R}$ in tokamak geometry, with $s=\sqrt{\psi/\psi_{\rm edge}}$ the normalized radial coordinate expressed in terms of the poloidal magnetic flux function $\psi$ and its value at the edge of the radial boundary $\psi_{\rm edge}$. ${\theta^{\star}}$ is the straight-field-line poloidal angle, and $\varphi$ is the toroidal angle. As this work is concerned with electrostatic turbulence, the equations of motion of $\vec{Z}$ are given by
@@
In order to reach quasi-steady state, it was shown Krommes (1999) that some form of dissipation is required. Given that we neglect physical collisions, we therefore introduce a noise control Krook operator $S_{\rm nj}=-\gamma_{\rm n}(f_{\rm j}-f_{\rm 0j})+S_{\rm nj}^{\rm c}$ of rate $\gamma_{\rm n}$, which is species independent and is taken to be a small fraction (typically around $10\%$) of the maximum linear growth rate of the instabilities driving the turbulence. $S_{\rm nj}$ relaxes the distribution $f_{\rm j}$ to the background $f_{\rm 0j}$, while conserving f.s.a. density, parallel moment, residual zonal flows, and energy. These conservation properties are ensured by the correction term $S_{\rm nj}^{\rm c}$. Besides serving as a noise control operator for the PIC scheme, temperature-gradient-driven simulations of this work use the source term $S_{\rm nj}$ as a fixed power heat source by imposing conservation of all f.s.a. moments except kinetic energy. For flux-driven simulations, i.e. with fixed prescribed power heat source, the f.s.a. heating operator used is given by
@@
Here, $\mathcal{E}=v^{2}/2=v_{\parallel}^{2}/2+\mu B$ is the kinetic energy per mass of the gyrocenter. The heat source $S_{\rm H}$ is parameterized by the f.s.a. profiles $n_{\rm H}$ and $T_{\rm H}$ appearing in Eq.(8), which in this work are taken to be the initial $n_{0}$ and $T_{0}$ profiles of each species. $\gamma_{\rm H}$ and $G_{\rm H}$ of Eq.(7) represent the heating rate and radial heating profile normalized to the unit of temperature, respectively. The form $\frac{\partial f_{\rm L}}{\partial T_{\rm H}}$ analytically ensures no f.s.a. density or momentum source. Nonetheless, $S_{\rm H}^{\rm c}$ ensures numerical conservation of f.s.a. density, parallel momentum, and residual zonal flows to round-off. Finally, to damp turbulence at the outer radial edge ($s=1$) of the simulation domain, a buffer in the form of a Krook operator $S_{\rm bj}(s)=-\gamma_{\rm b}[(s-s_{\rm b})/(1-s_{b})]^{4}\delta f_{j}$, of species-independent damping rate $\gamma_{\rm b}$ is used. $S_{\rm b}$ is parametrized by the species-dependent buffer entrance $s=s_{\rm b}$ and only acts in the radial range $s\in[s_{\rm b},1]$. Taken together, the r.h.s. of Eq.(1) now reads
@@ III Numerical methods
Under the delta-f splitting defined by Eq.(4), we choose $f_{0}=f_{\rm eq}$, where $f_{\rm eq}$ is an equilibrium distribution of the unperturbed system in the absence of the heat source term $S_{\rm H}$. This implies that $f_{\rm eq}$ has to be a function of the constants of motion of the unperturbed system. Namely, the modified canonical toroidal momentum Angelino et al. (2006) ${\hat{\psi}_{0}}=\psi+F(\psi)v_{\parallel}/\Omega_{\rm c}+\psi_{0corr}$, the kinetic energy $\mathcal{E}$, and the magnetic moment $\mu$. Here, $F(\psi)$ is the poloidal current flux function, as it appears in the representation of the axisymmetric equilibrium magnetic field $\vec{B}=F(\psi)\nabla\varphi+\nabla\psi\times\nabla\varphi$, and $\psi_{0corr}$ is a correction term that results in ${\hat{\psi}_{0}}\sim\psi$ on average along unperturbed trajectories. Under the Monte-Carlo interpretation Aydemir (1994), $f_{0}$ serves as a control variate, provided that the delta-f assumption
@@
Note that time dependence enters the control variate of Eq.(LABEL:eq:f0lm) via the f.s.a. background profiles of density $n_{0}$, parallel flow $u_{0}$ and temperature $T_{0}$ of the gyrocenters. In the following, Eq.(LABEL:eq:f0lm) will be used to develop the time dependent control variate scheme.
@@
must be appended. Given the analytic form for $f_{0}$ from Eq.(LABEL:eq:f0lm), Eq.(LABEL:eq:qne_corr) is calculated using quadratures on a 5D phase space grid. An alternative, and computationally cheaper way to avoid appending the correction term is to equate the two terms of Eq.(LABEL:eq:qne_corr). This means that the change in the background electron density is set to be approximately equal to the change in the background ion gyrodensity, i.e.
@@ IV Profiles and simulation parameters
where $\rho_{V}$ is the radial coordinate $\rho_{V}(\psi)=\sqrt{V(\psi)/V(\psi_{\rm edge})}$, with $V(\psi)$ the volume enclosed by the flux surface label $\psi$. $g_{0}$ and $g_{2}$ are coefficients determined such that $G$ and $\,\text{d}G/\,\text{d}\rho_{V}$ are continuous at $\rho_{V}=\rho_{\rm core}$, and $G_{\rm ped}=G_{1}+\mu_{G}(1-\rho_{\rm ped})$. The functional form, was found to describe well the experimentally observed profiles in TCV L-mode discharges Sauter et al. (2014). The values of the parameters in Eq.(LABEL:eq:profile) for the ion and electron density and temperature profiles are shown in Tab. 1. All densities and temperatures are normalized by $\bar{n}=N_{\rm ph}/V(\psi_{\rm edge})$ and $T_{0e}(s_{0})$ respectively. Here, $N_{\rm ph}$ is the physical total number of particles.
@@
The grid resolution for the radial $s$, poloidal ${\theta^{\star}}$ and toroidal $\varphi$ is taken to be $N_{s}\times N_{{\theta^{\star}}}\times N_{\varphi}=256\times 512\times 256$, where $N$ represents the number of intervals. Toroidal Fourier modes in the range of $0\leq n\leq 64$ will be simulated, only keeping poloidal Fourier modes $m\in[nq(s)-\Delta m,nq(s)+\Delta m]$ with half-width $\Delta m=5$. To lighten the computational cost ‘heavy’ electrons are considered, i.e. with an electron-ion mass ratio $m_{e}/m_{i}=1/200$. The time resolution is $c_{s}\Delta t/a=4.1\times 10^{-3}$, and the maximum linear growth rate is found to be $\gamma_{\rm max}a/c_{s}=0.931$. The strength of the Krook operator $S_{\rm n}$ for noise control is set to $\gamma_{\rm n}=11\%\gamma_{\rm max}$. For the flux-driven simulations considered here, the radial heat source profile $G_{\rm H}(\psi)$ of Eq.(7) is approximated by fitting a Gaussian function around the peak of the time-averaged effective heat source at quasi-steady state of a previously run temperature-gradient-driven simulation with the initial profiles given by Tab. 1. The heat source strength $\gamma_{\rm H}$ is determined by equating the integrated power to that of the effective heat source. The heat sink at the radial edge is replaced by the Krook buffer $S_{b}$. The calculation for the heat source is done for both ions and electrons separately. An example of the heat source profiles for ions and electrons are shown in Fig.2.
Simulations are initialized by loading markers in phase space and initializing their weights using quasi-random distributions. All adaptive cases have adaptation rates $\alpha_{n}=\alpha_{E}=11\%\gamma_{\rm max}$ and adaptation period set with $N_{t}=100$ for both ions and electrons. Even when considering the hybrid electron model, adapted electron profiles include contributions from both passing and trapped electrons. Unless stated otherwise, all time traces have a moving time averaging window of $c_{s}\Delta t/a=4.085$. Marker numbers $N_{p}$ are displayed in millions (M). Comparison will be made between the non-adaptive (standard) case with $N_{p}=256$M, $128$M and the adaptive case with $N_{p}=128$M, $64$M. The choice of simulations with different $N_{p}$ allows for the discerning of sampling noise’s effect on different diagnostic measures.
(a) Ions
(b) Electrons
@@ V.1 Time traces
We define an effective heat diffusivity $\chi$ from the f.s.a. radial component of the heat flux $\vec{q}_{H}$ resulting from gyrocenter $E\times B$ drifts. $\vec{q}_{H}$ is composed of the contributions from the kinetic energy $\vec{q}_{\rm kin}$, the field potential $\vec{q}_{\rm pot}$, and the particle flux $\vec{\Gamma}$, all of which are given respectively by Vernay et al. (2010)
@@
The heat effective diffusivity $\chi$ is then given by
@@
where $n$ and $T$ are the gyrocenter density and temperature respectively. The diffusivity is usually given in gyroBohm units $\chi_{GB0}=\rho_{s}^{2}c_{s}/a$ evaluated at $s=s_{0}=1$. Figure 3 shows the ion heat diffusivity $\chi_{i}$ radially averaged over $s\in[0.7,0.9]$ for all cases. Within the standard (non-adaptive) and adaptive cases, results appear converged in marker number $N_{p}$. The ion heat $\chi_{i}$ values of the adaptive cases are shown to be consistently higher than the standard cases. This difference between results of the two schemes can be explained by considering the shearing rate. Figure 4 shows the $s\in[0.7,0.9]$ radially averaged absolute value of the turbulence driven $E\times B$ zonal flow shearing rate, which is defined by Hahm (1994)
@@
To see how electron diffusivity $\chi_{e}$ compares to $\chi_{i}$, Fig.3(b) shows the ratio $\chi_{i}/\chi_{e}$ for all cases considered. In the time interval $c_{s}t/a\in[0,150]$, the system evolves gradually from an initially TEM-dominated regime to a mixed TEM-ITG regime, with a $\chi_{i}/\chi_{e}$ ratio evolving from about $0.2$ to $1.2$. For further times, this ratio remains at the same value. This change of regime is due essentially to the density profile evolution, as will be discussed further (see Fig.11).
@@
(a) Standard, ions
(b) Standard, electrons
(c) Adaptive, ions
(d) Adaptive, electrons
As the system approaches quasi-steady state, with the buffer being the only source/sink of particles, the particle flux is expected to reduce and reach near zero values. Figure 5 shows the f.s.a. ion gyrocenter flux $\left\langle\vec{\Gamma}\cdot\nabla\psi\right\rangle/\left\langle|\nabla\psi|\right\rangle$ for cases with various $N_{p}$ under standard and adaptive schemes. We first note that adaptive cases (blue and green curves) have consistent slightly higher ion gyrocenter flux compared to the standard cases (black and orange) right after the initial overshoot in the time window $c_{s}t/a\in[50,200]$. This is again consistent with the lower $\omega_{E\times B}$ amplitudes of the adaptive cases (see Fig.4), thereby allowing for higher turbulence levels, and therefore, particle transport to develop. Nonetheless, the standard case with $N_{p}=256$M (black line) seems to initially converge to the same quasi-steady state value as that of the adaptive cases.
@@
with $\left\langle w\right\rangle(s)=\sum_{p\in{\rm bin}}w_{p}/N_{\rm bin}$ and $\left\langle w^{2}\right\rangle(s)=\sum_{p\in{\rm bin}}w_{p}^{2}/N_{\rm bin}$ the expectation value of $w$ and $w^{2}$ within that bin, respectively. Here, $N_{\rm bin}$ is the number of markers belonging to the bin. Note that, one expects convergence of $\sigma_{w}$ with marker number $N_{p}$ at a given time $t$. Furthermore, with dissipation in the form of the Krook operator $S_{n}$, the growth of $\sigma_{w}$ will be limited at quasi-steady state. Figure 6 shows $\sigma_{w}$ for both standard and adaptive cases with marker number $N_{p}=128$M, up to a simulation time of $c_{s}t/a=600$. From Figs.6(a) and 6(b), we see that the radial location of increasing maximum values of $\sigma_{w}$ for both the ions and electrons are located where the profile deviation is the greatest (see Figs.10 and 11). After an initial phase where $\sigma_{w}$ grows (however to significantly lower values than in standard cases) the corresponding results under the adaptive scheme, on the other hand, have diminishing maxima. This is due to the adaptive control variate $f_{0}(t)$, with adaptive density and temperature profiles. Furthermore for the adaptive cases, we see a drop in electron $\sigma_{w}$ around $c_{s}t/a=50$, compared to that of the ion which drops gradually after $c_{s}t/a\simeq 200$, where $\sigma_{w}$ is gradually decreased. This can be explained by considering Fig.3(b), where we see that for $c_{s}t/a\in[0,50]$, the electron diffusivity $\chi_{e}$ is more than $200\%$ that of the ion $\chi_{i}$. This suggests a quicker evolving electron temperature profile $T_{e}$ during this period. Thus, as the adaptation rate is sufficiently high for this case, the electron $\sigma_{w}$ is reduced at a shorter time scale. Comparing results between the standard and adaptive cases, we see that the latter only have $15\%$ the maximum value of $\sigma_{w}$ compared to that of the former. Furthermore, at $c_{s}t/a=600$, $\sigma_{w}$ for the standard cases is still growing, while that of the adaptive cases exhibit a steady value of around $0.2$.
@@
Figure 8: Evolution up to time $c_{s}t/a=123$ of the radial profile of the zonal flow $E\times B$ shearing rate $\omega_{E\times B}$ with marker number $N_{p}=128$M. (a) Standard (non-adaptive). Adaptive cases (b) including correction term (LABEL:eq:qne_corr) calculated via quadratures, (c) imposing Eq.(LABEL:eq:qnrhs_fsa) and (d) neglecting correction term but still evolving $f_{0i}$ and $f_{0e}$ without imposing Eq.(LABEL:eq:qnrhs_fsa). All plots share the same color scale.